# A Robust Mixed-Effects Bandit Algorithm for Assessing Mobile Health Interventions

### v3

@@ 4 Algorithm Components
We now construct $\tilde{R}_{i,t}^{f}$, introduce the weighted least-squares loss function used by DML-TS-NNR to estimate parameters, and use nearest neighbor regularization to pool information across users.
@@ 4.1 Debiased Machine Learning
To construct $\tilde{R}_{i,t}^{f}$, we fit a working model $f_{i,t}$ for the conditional mean reward via the DML framework, which requires three main conditions: (a) Neyman Orthogonality - making the pseudo-outcome $\tilde{R}_{i,t}^{f}$ robust to misspecification of $f_{i,t}$, (b) sufficiently fast convergence of the working conditional mean model $f_{i,t}$, and (c) independence of $f_{i,t}$ and the observed outcomes. (a) is ensured by the structure of the doubly-robust differential reward. (b) requires us to introduce the following assumption:
@@
The working model, $f_{i,t}$, satisfies $\mathbb{E}_{p(s,\bar{a})}\left[(r_{i,t}(s,\bar{a})-f_{i,t}(s,\bar{a}))^{2}\right]=o_{P}(k^{-1/2})$ and $\mathbb{E}_{p(s)}\left[(r_{i,t}(s,0)-f_{i,t}(s,0))^{2}\right]=o_{P}(k^{-1/2})$.
Examples of supervised learning methods that converge at this rate (or faster) include parametric models of fixed dimension, local mean smoothers with small dimension (Wasserman, 2006, Chapter 5), and random forests and feedforward neural networks under certain regularity conditions (Biau, 2012; Farrell et al., 2021). Because of the Neyman Orthogonality, we can achieve sub-linear regret even if $f_{i,t}$ is misspecified. Assumption 3 guarantees better asymptotic variance than action-centering (AC) bandits Greenewald et al. (2017). Without it, the worst case performance (i.e., extreme model misspecification) can lead infinite asymptotic variance and perform empirically much worse than AC bandits (Bang & Robins, 2005). In mHealth contexts, we believe the assumption is fairly weak and plausible. The final requirement of independence, (c), can be accomplished via sample-splitting techniques. Below, we present two such options.
Option 1: The first option exploits the fact that outcomes across different users are independent. Consequently, we can perform sample splitting as follows. Step 1: Randomly assign each user $i$ to one of $J$ folds. Let $I_{j}$ denote the set containing user indexes for the $j$-th fold and let $I_{j}^{\complement}$ denote its complement. Step 2: For each fold, use a supervised learning algorithm to estimate the working model for $r_{i,t}(s,a)$ denoted $\hat{f}_{i,t}^{(j)}(s,a)$ using $I_{j}^{\complement}$. Step 3: Construct the pseudo-outcomes using (3).
Option 2: Because we do not expect an adversarial environment in mHealth, we may be willing to assume that the errors, $\epsilon_{i,t}$, are independent across time. In this case, we can adapt the procedure given above by randomly assigning data to folds at the level of $(i,t)$. In this case, $I_{j}$, contains the $(i,t)$ pairs assigned to fold $j$. This option is more powerful than option 1 in that it enables us to learn heterogeneity across users in the nuisance function, $f_{i,t}$, leading to greater variance reduction.
Lastly, we note two strategies that can reduce the computational demands of sample splitting. The first is to employ certain stable estimators that do not require sample splitting, e.g., bagged estimators using subsampling, as discussed in Chen et al. (2022). The second is to fit $f_{i,t}$ in an online fashion.
@@ 4.2 Estimation via Weighted Least Squares
Having formed $\tilde{R}_{i,t}^{f}(s,\bar{a})$, we now explain how to estimate the parameters, $\theta_{i,t}\in\mathbb{R}^{p}$, determining the differential reward, $\Delta_{i,t}(s,a)$. We assume that $\theta_{i,t}$ consists of three components:
@@
The parameter $\theta^{\text{shared}}$ is a “structural parameter” shared across all $i,t$. In contrast, $\theta^{\text{user}}_{i}$ and $\theta^{\text{time}}_{t}$ are “incidental parameters” specific to a given user and time point, respectively. We place these parameters in a single column vector, $\theta$, as $\textrm{vec}(\theta^{\text{shared}},\theta^{\text{user}}_{1},\ldots,\theta^{\text{user}}_{K},\theta^{\text{time}}_{1},\ldots,\theta^{\text{time}}_{K})$, where $K$ is the total number of stages. We then form a corresponding feature vector, $\phi_{i,t}=\phi(x_{i,t})$, where $\phi$ inserts $x_{i,t}$ into the positions corresponding to $\theta^{\text{shared}}$, $\theta^{\text{user}}_{i}$, and $\theta^{\text{time}}_{t}$ with zeros in all other locations; i.e., letting $p:=\text{dim}(x_{i,t})$, and the feature vector $\phi(x_{i,t})\in\mathbb{R}^{(2K+1)p}$ is defined as follows:
@@
We could then estimate $\theta_{i,t}$ via a weighted least squares (WLS) regression that minimizes the following objective function:
$\ell_{\text{WLS}}(\theta)=\sum_{i=1}^{n}\sum_{t=1}^{K-i+1}\tilde{\sigma}_{i,t}^{2}\big{(}\tilde{R}_{i,t}^{f}-\phi_{i,t}^{\top}\theta\big{)}^{2},$ (5)
where $\tilde{\sigma}_{i,t}^{2}=\pi_{i,t}(0|S_{i,t})\cdot(1-\pi_{i,t}(0|S_{i,t}))$. However, $\ell_{\text{WLS}}(\theta)$ is over-parameterized and does not leverage the available network information. The next section modifies $\ell_{\text{WLS}}(\theta)$ by adding regularization, producing the final loss function used in DML-TS-NNR.
@@ 4.3 Nearest-Neighbor Regularization
We assume access to network information relating neighboring users and time points. These networks are defined by graphs $G_{\text{user}}=(V_{\text{user}},E_{\text{user}})$ and $G_{\text{time}}=(V_{\text{time}},E_{\text{time}})$, where $V$ denotes vertices and $E$, edges. In practice, these networks can be formed via known clusters (e.g., shared schools, companies, geographic locations) or similar covariates. In the case of $G_{\text{time}}$, a reasonable choice is to construct a graph of neighboring sequential time points; i.e., $t=1\leftrightarrow t=2\leftrightarrow t=3\dots$. We assume that neighboring units have similar incidental parameters as follows:
Assumption 5.
There exists $D_{\text{user}},D_{\text{time}}\in\mathbb{R}^{+}$ such that $\|\theta^{\text{user}}_{i}-\theta^{\text{user}}_{j}\|_{2}^{2}\leq D_{\text{user}}$ and $\|\theta^{\text{time}}_{i}-\theta^{\text{time}}_{j}\|_{2}^{2}\leq D_{\text{time}}$, for all $i,j$.
Assumption 5 suggests we may be able to improve our estimates of $\theta$ by regularizing neighboring values of $\theta^{\text{user}}_{i}$ and $\theta^{\text{time}}_{i}$ toward each other. This can be accomplished by using an $L^{2}$ Laplacian penalty. We illustrate the idea using $G_{\text{user}}$; the application to $G_{\text{time}}$ is similar.
First, we form the incidence matrix, $Q\in\mathbb{R}^{N\times N}$, with the entry $Q_{v,e}$ corresponding to the $v$-th vertex (user) and $e$-th edge. Denote the vertices of $e$ as $v_{i}$ and $v_{j}$ with $i>j$. $Q_{v,e}$ is then equal to 1 if $v=v_{i}$, -1 if $v=v_{j}$, and 0 otherwise. The Laplacian matrix is then defined as $L=QQ^{\top}$. Similar to Yang et al. (2020), we can then form a network cohesion penalty across users as follows:
@@
where $\Theta_{\text{user}}:=(\theta^{\text{user}}_{1},\ldots,\theta^{\text{user}}_{N})^{\top}\in\mathbb{R}^{N\times p}$ and $L_{\text{user}}$ is the Laplacian matrix for users. The penalty is small when $\theta^{\text{user}}_{i}$ and $\theta^{\text{user}}_{j}$ are close for connected users. We employ a shared regularization hyperparameter $\lambda$, for $G_{\text{user}}$ and $G_{\text{time}}$, and include a standard $L^{2}$ penalty on $\theta$ with hyperparameter $\gamma$. The full penalization matrix $V_{0}\in\mathbb{R}^{(2K+1)p\times(2K+1)p}$ is
$V_{0}=\operatorname{diag}(\gamma I_{p},\lambda L^{\text{user}}_{\otimes}+\gamma I_{Kp},\lambda L^{\text{time}}_{\otimes}+\gamma I_{Kp}),$ (6)
where $L^{\text{user}}_{\otimes}:=L_{\text{user}}\otimes I_{p}$ , the Kronecker product of $L_{\text{user}}$ with $I_{p}$. We then adapt $\ell_{\text{WLS}}(\theta)$ to include a corresponding penalty term:
@@
In Section 5 we show that (7) leads to a Thompson sampling algorithm for action selection.
@@ 5.1 Algorithm Description
The minimizer of the weighted regularized least squares loss in (7) can be expressed as $\hat{\theta}=V^{-1}b$, where
$V=V_{0}+\sum_{i=1}^{n}\sum_{t=1}^{K-i+1}\tilde{\sigma}_{i,t}^{2}\phi_{i,t}\phi_{i,t}^{\top},\quad b=\sum_{i=1}^{n}\sum_{t=1}^{K-i+1}\tilde{R}_{i,t}^{f}\sigma_{i,t}^{2}\phi_{i,t}.$
In analogy to Bayesian linear regression, the penalized and weighted Gram matrix, $V$, plays the role of a posterior precision matrix. This motivates an algorithm in which we randomly perturb the point estimate, $\hat{\theta}$, scaling by an inverse square-root matrix, $V^{-1/2}$, times a mean-zero random deviate, $\eta$. The perturbed estimate, $\tilde{\theta}$, can be written as $\tilde{\theta}=\hat{\theta}+\beta(\delta)V^{-1/2}\eta,$ where $\beta(\delta)$ is defined as
$\beta(\delta)$ $=v\left[2\log\left(\frac{\det(V)^{1/2}}{\det(V)^{1/2}\delta/2}\right)\right]^{1/2}+B,$ (8)
$B$ $=K\frac{\lambda}{\sqrt{\gamma}}\left(D_{\text{user}}+D_{\text{time}}\right)+\sqrt{\gamma n}\left(B_{\text{user}}+B_{\text{time}}\right).$ (9)
The value $v$ is defined in Corollary 4 and can be set to a sufficiently large constant. The hyperparameter $\delta$ determines the probability with which the regret bound holds. The random deviate, $\eta$, is drawn from a distribution, $\mathcal{D}^{TS}$, satisfying the technical conditions given in Definition 1 in Appendix E. In practice, we recommend setting $\mathcal{D}^{TS}$ to a multivariate Gaussian distribution or t-distribution to simplify calculations. Given $\tilde{\theta}$, we then compute $\bar{A}_{i,t}=\operatorname{argmax}_{a\in[q]\backslash\{0\}}\phi_{i,t}^{\top}\tilde{\theta}$ and randomly select $A_{i,t}=0$ with probability
$\pi_{i,t}(0|\bar{A}_{i,t}):=\max\big{(}\pi_{\min},\min\{\pi_{\max},\Pr(\phi_{i,t}^{\top}\tilde{\theta}>0)\}\big{)},$ (10)
@@
  $\tilde{\theta}=\hat{\theta}+\beta(\delta)V^{-1/2}\eta$, with $\beta(\delta)$ given in (8)
  $\bar{A}_{i,t}=\operatorname{argmax}_{a\in[q]\backslash\{0\}}\phi_{i,t}^{\top}\tilde{\theta}$
@@
Having detailed the action selection algorithm, we now summarize DML-TS-NNR in Algorithm 2. The algorithm includes an outer loop that iterates through the stages from $k=1$ to $k=K$. Within each stage, we handle users sequentially, iteratively selecting actions according to Algorithm 1, observing rewards, and updating $V$ and $b$ accordingly.
Algorithm 2 DML-TS-NNR
@@
        Observe rewards, $R_{i,t}$
        Construct pseudo-rewards, $\tilde{R}_{i,t}^{f^{(m)}}$, as explained in Section 4.1
        Set $\phi=\phi(x_{i,t})$ as described in (4)
@@ 5.2 Regret bound
$\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\big{[}\pi_{i,t}^{\star}(\bar{A}_{i,t}^{*}|S_{i,t})\cdot x(S_{i,t},\bar{A}_{i,t}^{*})^{\top}\theta^{*}_{i,t}\\ -\pi_{i,t}(\bar{A}_{i,t}|S_{i,t})\cdot x(S_{i,t},\bar{A}_{i,t})^{\top}\theta^{*}_{i,t}\big{]}.$
The inner sum calculates regret (the difference between the expected reward of the optimal action and the chosen action) weighted by the probability of the action, following Greenewald et al. (2017), for all individuals at stage $k$. We then average these values to account for the growing number of pairs $(i,t)$ as the stages progress. The cumulative regret is the sum of these averaged values over all stages.
The rewards associated with the baseline action are excluded since $\Delta_{i,t}(S_{i,t},0)=0$. $\boldsymbol{\operatorname{Regret}}_{K}$ counts as a pseudo-regret because the randomness attributed to $\{\epsilon_{i,t}\}_{t=1}^{K}$ has been eliminated (Audibert et al., 2003). Before presenting the regret bound, we introduce one more assumption:
@@
$\|x(s,a)\|\leq 1$ for all $s$ and $a$. Also, there exists a known value $M\in\mathbb{R}^{+}$ bounding both $\|\theta_{i,t}\|$ and $|g_{t}(s)|$ for all $i$, $t$, and $s$. Further, $|f_{i,t}|\leq 2M$.
@@
Under Assumptions 16, with probability at least $1-\delta$, $\boldsymbol{\operatorname{Regret}}_{K}$ satisfies
$O\left(\left(\beta_{K}(\delta^{\prime})+\gamma_{K}(\delta^{\prime})\left[1+\frac{4}{p}\right]\right)\boldsymbol{\cdot}\log(K)\sqrt{4cKp}\right)$
where $\delta^{\prime}=\delta/4K$ and $\min(\pi(0|s),1-\pi(0|s))>1/c$. The exact regret bound is given and both $\beta_{K}$ and $\gamma_{K}$ are defined in Appendix F.
Theorem 1’s proof is in Appendix F. It resembles  Abeille & Lazaric (2017), albeit with adjustments to address (a) the nonlinear baseline and (b) the increasing pool of users and time points. The non-linear baseline requires deriving the new sub-Gaussian variance factor for the centered pseudo-reward and adapting classic contextual bandit theory (Abbasi-Yadkori et al., 2011) to our weighted least squares setting. The increasing pool of users and time points requires modifying the Abeille & Lazaric (2017) proof with a careful use of stages and Cauchy Schwarz. The overall bound scales as
$O\left(p^{3/2}\sqrt{\log^{2}(K)K}\right).$
DML is particularly advantageous for $\beta_{K}(\delta^{\prime})$ and $\gamma_{K}(\delta^{\prime})$, which rely on the convergence rate of the model $f$ towards the true mean differential reward $r$, as discussed further in Appendix F.2. Another interesting aspect of the regret bound is its dependence solely on the dimension, $p$, of the differential reward model, rather than the dimension of the baseline reward. This characteristic allows us to achieve robust regret bounds even in cases where the baseline reward is highly complex.

### v4

@@ 4 Algorithm Components
We now construct $\tilde{R}_{i,t}^{f}$, introduce the weighted least-squares loss function used by RoME to estimate parameters, and use nearest-neighbor regularization to pool information across users.
@@ 4.1 Debiased Machine Learning
To construct $\tilde{R}_{i,t}^{f}$, we fit a working model $f_{i,t}$ for the conditional mean reward. Because $\tilde{R}_{i,t}^{f}$ is robust to misspecification of $f_{i,t}$, our regret bound in Theorem 1 does not require $f_{i,t}$ to be correctly specified. However, we show in Appendix E.2 that the following assumption results in a lower asymptotic variance bound for $\tilde{R}_{i,t}^{f}$.
@@
The working model, $f_{i,t}$, satisfies $|r_{i,t}(s,a)-f_{i,t}(s,a)|\overset{a.s.}{\to}0$ for all $s$, $a$.
Assumption 3 states that $f_{i,t}$ is a strongly consistent estimator of $r_{i,t}$, but it does not impose any specific rate requirement on the convergence. Although our regret bound does not require Assumption 3, it does require two standard conditions within the DML framework: (a) Neyman orthogonality and (b) independence of $f_{i,t}$ and the observed rewards. The construction of $\tilde{R}_{i,t}^{f}$ ensures that (a) holds, making $\tilde{R}_{i,t}^{f}$ robust to misspecification of $f$. We can enforce (b) via sample-splitting techniques. We present two such options below.
Option 1: The first option exploits the fact that outcomes across different users are independent. Consequently, we can perform sample splitting as follows. Step 1: Randomly assign each user $i$ to one of $J$ folds. Let $W_{j}$ denote the set containing user indexes for the $j$-th fold and let $W_{j}^{\complement}$ denote its complement. Step 2: For each fold, use a supervised learning algorithm to estimate the working model for $r_{i,t}(s,a)$ denoted $\hat{f}_{i,t}^{(j)}(s,a)$ using $W_{j}^{\complement}$. Step 3: Construct the pseudo-outcomes using (1).
Option 2: Because we do not expect an adversarial environment in mHealth, we may be willing to assume that the errors, $\epsilon_{i,t}$, are independent across time. In this case, we can adapt the procedure given above by randomly assigning data to folds at the level of $(i,t)$. In this case, $W_{j}$, contains the $(i,t)$ pairs assigned to fold $j$. This option is more powerful than option 1 in that it enables us to learn heterogeneity across users in the nuisance function, $f_{i,t}$, leading to greater variance reduction.
Lastly, we note two strategies that can reduce the computational demands of sample splitting. The first is to employ estimators that satisfy a leave-one-out stability condition, such as bagged estimators that use subsampling (Chen et al., 2022); these estimators eliminate the need to perform sample splitting. The second is to fit $f_{i,t}$ in an online fashion.
@@ 4.2 Estimation via Weighted Least Squares
Having formed $\tilde{R}_{i,t}^{f}(s,\bar{a})$, we now explain how to estimate the parameters, $\theta_{i,t}\in\mathbb{R}^{d}$, determining the differential reward, $\Delta_{i,t}(s,a)$. We assume that $\theta_{i,t}$ consists of three components:
@@
The parameter $\theta^{\text{shared}}$ is a fixed effect shared across all $i,t$. In contrast, $\theta^{\text{user}}_{i}$ and $\theta^{\text{time}}_{t}$ are random effects specific to a given user and time point, respectively. We place these parameters in a single column vector, $\theta$, as $\textrm{vec}(\theta^{\text{shared}},\theta^{\text{user}}_{1},\ldots,\theta^{\text{user}}_{K},\theta^{\text{time}}_{1},\ldots,\theta^{\text{time}}_{K})$, where $K$ is the total number of stages. We then form a corresponding feature vector, $\phi_{i,t}=\phi(x_{i,t})\in\mathbb{R}^{(2K+1)d}$, where $\phi$ inserts $x_{i,t}$ into the positions corresponding to $\theta^{\text{shared}}$, $\theta^{\text{user}}_{i}$, and $\theta^{\text{time}}_{t}$ with zeros in all other locations; i.e., $\phi(x_{i,t})=C_{i,t}^{{\top}}x_{i,t}$, where
@@
so that $C_{i,t}\in\mathbb{R}^{d\times(2K+1)d}$. We could then estimate $\theta_{i,t}$ via a weighted least squares (WLS) regression that minimizes the following objective function:
$\ell_{\text{WLS}}(\theta)=\sum_{i=1}^{n}\sum_{t=1}^{K-i+1}\tilde{\sigma}_{i,t}^{2}\big{(}\tilde{R}_{i,t}^{f}-\phi_{i,t}^{\top}\theta\big{)}^{2},$
where $\tilde{\sigma}_{i,t}^{2}=\pi_{i,t}(0|S_{i,t})\cdot\{1-\pi_{i,t}(0|S_{i,t})\}$. However, $\ell_{\text{WLS}}(\theta)$ is over-parameterized and does not take advantage of available network information. The next section modifies $\ell_{\text{WLS}}(\theta)$ by adding regularization, producing the final loss function used in RoME.
@@ 4.3 Nearest-Neighbor Regularization
We assume access to network information relating neighboring users and time points. These networks are defined by graphs $G_{\text{user}}=(V_{\text{user}},E_{\text{user}})$ and $G_{\text{time}}=(V_{\text{time}},E_{\text{time}})$, where $V$ denotes vertices and $E$, edges. In practice, these networks can be formed via known clusters (e.g., shared schools, companies, geographic locations) or similar covariates. In the case of $G_{\text{time}}$, a reasonable choice is to construct a graph of neighboring sequential time points; i.e., $t=1\leftrightarrow t=2\leftrightarrow t=3\dots$. Because we expect neighboring users and time points to have similar parameters, we may be able to improve our estimates of $\theta$ by regularizing neighboring values of $\theta^{\text{user}}_{i}$ and $\theta^{\text{time}}_{t}$ toward each other. This can be accomplished by using an $L^{2}$ Laplacian penalty. We illustrate the idea using $G_{\text{user}}$; the application to $G_{\text{time}}$ is similar.
First, we form the incidence matrix, $Q$, with the entry $Q_{v,e}$ corresponding to the $v$-th vertex (user) and the $e$-th edge. Denote the vertices of this edge as $v_{i}$ and $v_{j}$ with $i>j$. $Q_{v,e}$ is then equal to 1 if $v=v_{i}$, -1 if $v=v_{j}$, and 0 otherwise. The Laplacian matrix is then defined as $L=QQ^{\top}\in\mathbb{R}^{K\times K}$. Similar to Yang et al. (2020), we form a network cohesion penalty across users as follows:
@@
where $\Theta_{\text{user}}\coloneq(\theta^{\text{user}}_{1},\ldots,\theta^{\text{user}}_{K})^{\top}\in\mathbb{R}^{K\times d}$ and $L_{\text{user}}$ is the Laplacian matrix for users. The penalty is small when $\theta^{\text{user}}_{i}$ and $\theta^{\text{user}}_{j}$ are close for connected users. We employ a shared regularization hyperparameter $\lambda$, for $G_{\text{user}}$ and $G_{\text{time}}$, and include a standard $L^{2}$ penalty on $\theta$ with hyperparameter $\gamma$. The full penalization matrix $V_{0}\in\mathbb{R}^{(2K+1)d\times(2K+1)d}$ is
$V_{0}=\operatorname{diag}(\gamma I_{d},\gamma I_{Kd}+\lambda L^{\text{user}}_{\otimes},\gamma I_{Kd}+\lambda L^{\text{time}}_{\otimes}),$ (2)
where $L^{\text{user}}_{\otimes}\coloneq L_{\text{user}}\otimes I_{d}$ , the Kronecker product of $L_{\text{user}}$ with $I_{d}$ ($L^{\text{time}}_{\otimes}$ is defined similarly). We then adapt $\ell_{\text{WLS}}(\theta)$ to include a corresponding penalty term:
@@
In Section 5, we show that (3) leads to a Thompson sampling algorithm for action selection.
@@ 5.1 Algorithm Description
The minimizer of the weighted regularized least squares loss in (3) is $\hat{\theta}=V^{-1}b$, where
$V=V_{0}+\sum_{i=1}^{n}\sum_{t=1}^{K-i+1}\tilde{\sigma}_{i,t}^{2}\phi_{i,t}\phi_{i,t}^{\top},\quad b=\sum_{i=1}^{n}\sum_{t=1}^{K-i+1}\tilde{R}_{i,t}^{f}\tilde{\sigma}_{i,t}^{2}\phi_{i,t}.$
[…]
$\beta_{i,t}(\delta)\coloneq v\sqrt{2\log\left\{\frac{2K(K+1)}{\delta}\right\}+\log\left\{\frac{\det(\underline{\hbox{\set@color$V$}}_{i,t}^{-1})}{\det(\Lambda_{i,t}^{0})}\right\}}+\zeta\max\{\log^{3/4}(K),1\}.$ (4)
In the above expression, $\Lambda_{i,t}^{0}\coloneq C_{i,t}V^{-1}V_{0}V^{-1}C_{i,t}^{{\top}}$ and $K$ is the total number of stages. The value $v$ is the sub-Gaussian factor for $\tilde{R}_{i,t}^{f}$, and $\zeta$ is defined in Remark 2 in Appendix E; both can be set to sufficiently large constants. The hyperparameter $\delta$ determines the probability with which the regret bound holds. The random deviate, $\eta$, is drawn from a distribution $\mathcal{D}^{TS}$ satisfying the technical conditions in Definition 1 in Appendix D. In practice, we recommend setting $\mathcal{D}^{TS}$ to a multivariate Gaussian distribution or t-distribution to simplify calculations. Given $\tilde{\theta}_{i,t}$, we then compute $\bar{A}_{i,t}=\operatorname{argmax}_{a\in[q]\backslash\{0\}}x_{i,t}^{\top}\tilde{\theta}_{i,t}$ and randomly select $A_{i,t}=0$ with probability
$\pi_{i,t}(0|\bar{A}_{i,t})\coloneq\max\left[\pi_{\min},\min\left\{\pi_{\max},\Pr(x_{i,t}^{\top}\tilde{\theta}_{i,t}<0)\right\}\right],$ (5)
@@
  Set $\tilde{\theta}_{i,t}=C_{i,t}\hat{\theta}+\beta_{i,t}(\delta)\underline{\hbox{\set@color$V$}}_{i,t}^{-1/2}\eta$
  Set $\bar{A}_{i,t}=\operatorname{argmax}_{a\in[q]\backslash\{0\}}x_{i,t}^{\top}\tilde{\theta}_{i,t}$
@@
Having detailed the action selection algorithm, we now summarize RoME in Algorithm 2. The algorithm includes an outer loop that iterates through the stages from $k=1$ to $k=K$. Within each stage, we handle users sequentially, iteratively selecting actions according to Algorithm 1, observing rewards, and updating $V$ and $b$ accordingly.
Algorithm 2 RoME
@@
        Observe reward, $R_{i,t}$
        Construct pseudo-reward, $\tilde{R}_{i,t}^{f}$, as explained in Section 4.1
        Set $\phi_{i,t}=C_{i,t}^{{\top}}x_{i,t}$
@@ 5.2 Regret bound
$\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\left\{\pi_{i,t}^{\star}(\bar{A}_{i,t}^{\star}|S_{i,t})\cdot x(S_{i,t},\bar{A}_{i,t}^{\star})^{\top}\theta^{\star}_{i,t}-\pi_{i,t}(\bar{A}_{i,t}|S_{i,t})\cdot x(S_{i,t},\bar{A}_{i,t})^{\top}\theta^{\star}_{i,t}\right\}.$
The inner sum calculates regret (the difference between the expected reward of the optimal action and the chosen action) weighted by the probability of the action, following Greenewald et al. (2017), for all users at stage $k$. We then average these values to account for the growing number of pairs $(i,t)$ as the stages progress. The cumulative regret is the sum of these averaged values over all stages.
The rewards associated with the baseline action are excluded since $\Delta_{i,t}(S_{i,t},0)=0$. $\boldsymbol{\operatorname{Regret}}_{K}$ is a form of pseudo-regret because it eliminates the randomness due to $\{\epsilon_{i,t}\}_{t=1}^{K}$ (Audibert et al., 2003). The regret bound requires additional technical assumptions:
@@
For all $s$ and $a$, $\|x(s,a)\|\leq 1$.
Assumption 6.
There exists a known value $M\in\mathbb{R}^{+}$ such that $|g_{t}(s)|,|f_{i,t}|\leq M$ for all $i$, $t$, and $s$.
Assumption 7.
Let $e_{K}\coloneq\#E^{\text{user}}_{K}+\#E^{\text{time}}_{K}$ denote the total number of edges. Then $e_{K}=O(K^{\omega})$ for some $\omega\in[0,\infty)$.
Assumption 8.
Let $\theta_{k}^{\star}$ represent the vector of true parameter values at stage $k=i+t-1$ and $V_{i,t}^{-1}$, the matrix $V$ prior to decision point $(i,t)$. Then $\sup_{(i,t)\in\mathcal{O}_{K}}\|V_{0}\theta_{k}^{\star}\|_{V_{i,t}^{-1}}=O_{p}\left\{\log^{3/4}(K)\right\}.$
Assumption 9.
[…]
Assumption 5 is a standard assumption that simplifies the proof (Abeille & Lazaric, 2017). Assumption 6 ensures that the (estimated) baseline rewards are bounded. Assumption 7 ensures that the number of edges grows no faster than a polynomial rate. Assumption 8 limits the contribution of the random effects to the regret bound. Lastly, Assumption 9 ensures sufficient exploration of the context space. See the proof of Theorem 1 in Appendix E.3 for additional discusion of Assumptions 8 and 9. We now state the regret bound.
@@
Under Assumptions 12, 49, with probability at least $1-\delta$, $\boldsymbol{\operatorname{Regret}}_{K}$ is of order
$O\left[d\sqrt{K\log(K)\log(K^{2}d)}\left\{\sqrt{d}+\log^{1/4}(K)\right\}\right].$
The exact bound (including constants) and proof are given in Appendix E. Following Greenewald et al. (2017), we decompose the regret into two terms. The first term depends on a summation of terms of the form $\{\pi_{i,t}^{\star}(\bar{A}_{i,t}^{\star}|S_{i,t})-\pi_{i,t}(\bar{A}_{i,t}|S_{i,t})\}x(S_{i,t},\bar{A}_{i,t})^{\top}\theta_{i,t}^{\star}$. We employ a novel technique based on Markov’s inequality to bound this term (see Lemmas 12 and 13 in Appendix E).
We bound the second term using the high-level proof technique of Abeille & Lazaric (2017). The primary difficulty in bounding this term is that the classical theory for the estimation error of RLS estimates (Theorem 2 in Abbasi-Yadkori et al. (2011)) is not directly applicable to our setting due to the increasing dimensionality. Lemma 6 in Appendix E generalizes this theory to our setting. This Lemma produces a bound involving $\|C_{i,t}V_{i,t}^{-1}V_{0}\theta_{k}^{\star}\|_{\underline{\hbox{\set@color$V$}}_{i,t}}$, which we subsequently bound by $\|V_{0}\theta_{k}^{\star}\|_{V_{i,t}^{-1}}$ in Lemma 7. This section of the proof requires Assumption 8 to ensure that the contribution of the new parameters to this bound is small enough that we incur no more than a logarithmic penalty.
The overall bound scales like $\tilde{O}(\sqrt{K}d^{3/2})$ up to log factors, matching the standard rate given in Agrawal & Goyal (2012) and Abeille & Lazaric (2017) for linear Thompson sampling with fixed dimension. In particular, the bound depends solely on the dimensionality of the differential reward—not the baseline reward, which may be much more complex.